OpenStudy (anonymous):

Which is the graph of the system of equations? x + 2y = 2 4x – y = 4

OpenStudy (anonymous):

@johnweldon1993 This is the question! I need a few moments to get the pictures though.

OpenStudy (johnweldon1993):

Alright, so, whats the first step? :)

OpenStudy (anonymous):

I believe the first step is to change the equations to Slope Intercept Form...

OpenStudy (johnweldon1993):

Indeed...so what is each equation in slope intercept form?

OpenStudy (anonymous):

Um, let me see...

OpenStudy (anonymous):

The first one is: x = 2 - 2y And the second one is: -y = 4 - 4x

OpenStudy (johnweldon1993):

Hmm, so remember how slope intercept looks like \[\large y = mx + b\] we want 'y' by itself in both cases...so try that 1 more time :)

OpenStudy (anonymous):

Oh, I thought it would be okay if it was x, but I will definitely try again. :)

OpenStudy (anonymous):

I think I got this incorrect, but: y = 1 - x

OpenStudy (johnweldon1993):

Not quite...but I'm glad you're trying :D So we'll do the first...first XD \[\large x + 2y = 2\] We want the 2y by itself...so lets subtract 'x' from both sides first \[\large 2y = -x + 2\] and now we want 'y' COMLETELY y itself...so divide everything by 2 \[\large y = -\frac{1}{2}x + 1\]

OpenStudy (anonymous):

I subtracted 'x' from both sides, but instead I put it like this: 2y = 2 - x

OpenStudy (anonymous):

So, where did you get the half from? That's what I don't understand...

OpenStudy (johnweldon1993):

Same thing there You can write it like that or like mine....same thing just different order :) So lets do it how you wrote to make it easier :) \[\large 2y = 2 - x\] Now...we want \(\large y = something\) so we need to make that 2 that is being multiplied to the 'y' go away...and how do we get rid of something being multiplied? We divide So what is 2y divided by 2? well it's just 1y...or just y So that means...if we divide everything by 2...we have \[\large \frac{\cancel{2}y = \cancel{2}\rightarrow 1 - x}{\cancel{2}} \rightarrow y = 1 - \frac{x}{2}\] which can then be written as \[\large y = 1 -\frac{1}{2}x\]

OpenStudy (johnweldon1993):

Just to address your exact question... what is x....divided by 2? well it is 1/2 'x' :)

OpenStudy (anonymous):

Ohhh! Okay, I think I understand better. So is the overall answer to the question D. ?

OpenStudy (anonymous):

@johnweldon1993 Am I right? Is the answer D. ?

OpenStudy (johnweldon1993):

oh sorry, looking at another question lol...yes D is correct :)

OpenStudy (anonymous):

It's okay, I took the dare and submitted in my answer! Turns out it was correct. :D Thanks so much for your help! :D